OpenStudy (mpj4):

How to find the 2nd derivative of 2x^2-3y^2-4?

OpenStudy (thomas5267):

You find the first derivative the differentiate the first derivative.

OpenStudy (mpj4):

Ah well is it 4x-6y dy/dx?

OpenStudy (anonymous):

That's the first derivative. Now differentiate it again ^^

OpenStudy (mpj4):

err I don't know how. Lemme try that again. Can you please explain how -3y^2 disappeared in this solution? This was the solution from a second derivative calculator

OpenStudy (mpj4):

It's the same equation. I didn't understand why -3y^2 was treated as a constant

OpenStudy (thomas5267):

Because the calculator is dumb.

OpenStudy (thomas5267):

I think you are not dealing with partial derivatives right?

OpenStudy (anonymous):

Me neither. It would have made sense if it was \[\Large \frac{\partial}{\partial x}\] though. I agree with Thomas, your calculator is dumb. Trust your human abilities :P

OpenStudy (anonymous):

Okay, do you know the product rule? ^^

OpenStudy (mpj4):

fg'+gf'?

OpenStudy (anonymous):

One thing to know it in theory, and another to practice it :P Let's do that right now. Time to differentiate \[\Large 4x - 6y \frac{dy}{dx}\] So... ready?

OpenStudy (mpj4):

ok, I'm ready

OpenStudy (anonymous):

Good. Time to differentiate. \[\Large \frac{d}{dx}\left[4x - 6y \frac{dy}{dx}\right]\] Let's start with the 'easy' bit. Differentiate 4x, you get?

OpenStudy (mpj4):

4

OpenStudy (anonymous):

Good :P \[\Large 4 - \frac{d}{dx}\left[6y\frac{dy}{dx}\right]\] Now you have the derivative of a product, 6y and dy/dx Differentiate.

OpenStudy (anonymous):

Hey, are you having trouble? Put down your best try, don't be afraid to make mistakes. OR to tell me that you're stuck. It's MUCH better that you get stuck/commit errors here than on an exam. ^_^

OpenStudy (mpj4):

Well first I took out the 6 and have y dy/dx left to derive

OpenStudy (mpj4):

then applying the product rule on y dy/dx I got stuck

OpenStudy (mpj4):

\[4-6(\frac{d}{dx}(y\frac{dy}{dx}))\]

OpenStudy (thomas5267):

\[ \frac{d}{dx}\left(\frac{dy}{dx}\right)=\frac{d^2y}{dx^2} \] Don't think too much.

OpenStudy (anonymous):

(fg)' = f'g + fg' It's all a matter of knowing exactly what your f and g are. Here's an idea. Your "f" here is y, and your "g" here is dy/dx

OpenStudy (mpj4):

\[y (\frac{d}{dx}\frac{dy}{dx})+\frac{dy}{dx}(\frac{d}{dx}y)\] like this?

OpenStudy (anonymous):

Yup. Pretty much.

OpenStudy (anonymous):

Be careful :)

OpenStudy (mpj4):

how am I supposed to do \[\frac{d}{dx}y\]

OpenStudy (anonymous):

Uhh... what indeed :P \[\Large \frac{d}{dx}y \qquad \rightarrow \qquad \frac{dy}{dx}\] don't you think? :P

OpenStudy (mpj4):

hmm so there's no special meaning behind it =____=

OpenStudy (anonymous):

nope. nothing more you can do about that.

OpenStudy (mpj4):

then I have \[4-6(\frac{d^2y}{dx^2}+(\frac{dy}{dx})^2)\]

OpenStudy (anonymous):

There you go ^^ That wasn't so hard, was it? :P

OpenStudy (mpj4):

turns out we were wrong ._.

OpenStudy (mpj4):

I failed the homework and showed me the solution

OpenStudy (anonymous):

We weren't. You just didn't show the entire problem :P That = 4 matters.

OpenStudy (anonymous):

Had we started with \[\Large 2x^2 - 3y^2 = 4\] We would have arrived there too ^^

OpenStudy (mpj4):

Oh well

OpenStudy (mpj4):

Thanks for the help :)

OpenStudy (anonymous):

Sorry we didn't arrive at your desired solution. Next time (if ever) show a screenshot of your problem ^^